MOLECULAR EMISSION STRUCTURE IN THE RED RECTANGLE - SEQUENCE BANDS?

\underline{TIMOTHY W. SCHMIDT}, NEIL J. REILLY, SCOTT H. KABLE, School of Chemistry, University of Sydney, NSW 2006, Australia; ROBERT G. SHARP, Anglo-Australian Observatory, PO Box 296 Epping, NSW 1710, Australia.

We report high resolution (R ≈ 37000) Integral Field spectroscopy of the central region (r < 14 arcsec) of the Red Rectangle nebula surrounding HD 44179. The observations focus on the 5800 Å emission feature, the bluest of the Red Rectangle Bands. We propose that the emission feature, widely believed to be a molecular emission band, is sequence structure emission from a molecule with an extended chromophore. We present model spectra for a range of PAH cations with elevated vibrational temperatures.